OpenStudy (anonymous):

-3 1/7 divided by -1 1/4?

OpenStudy (anonymous):

Convert to improper fraction, invert and multiply \[-22/7 \div -5/4 = -22/7 * -4/5 = 88/35\]

hero (hero):

Oops

hero (hero):

thought it was -31/7

OpenStudy (anonymous):

lol thats alright i still gave u a medal. thanks